OpenStudy (anonymous):

Write a slope-intercept equation for a line parallel to the line x + 7y = 3 which passes through the point (–14, 6).

OpenStudy (anonymous):

if it is parralel it will have the same slope so first we find the slope of the original line x+7y=3 so we would rearrange this to \[(3-x)/7\]

OpenStudy (anonymous):

so the equation would be\[(3/7)-(x/7)=y \] and the slope would be -(1/7)x

OpenStudy (anonymous):

so now using point slope form we can derive the new equation\[x-x _{1} =m(y-y _{1})\]